Startind from the origin a body oscilates simple harmonic with a period of 2s after what time will it`s kE energy be75% of total energy

Grade:11

1 Answers

DR STRANGE
7 years ago
Hi,
     K.E = (Total energy)*(1-x2) …......(considering amplitude to be 1)
where x is distance from origin
K.E. = 75%Total energy = ¾ total energy
1-x2=3/4  
x=1/2
now , x = \sin \omega t   …...(oscillation starts from origin)
And
\omega =\frac{2\pi }{T}
Given T=2sec
\frac{1}{2}=\sin \frac{2\pi }{T}t=\sin \frac{2\pi }{2}t=\sin \pi t
 Hence t= 1/6 sec... it will take 1/6 sec.
Please approve if you are satisfied :)